OpenStudy (anonymous):

Solve the equation x-5/4 = 3 for x. Please show your work and solution in the space provided.

OpenStudy (anonymous):

transpose the \[-\frac{ 5 }{ 4 }\] to the other side then change its sign then perform the necesssary operation then thats it \[x=\frac{ 17 }{ 4 }\] check \[\frac{ 17 }{ 4 }-\frac{ 5 }{ 4 }=3\]

OpenStudy (anonymous):

i have no idea what that says...

OpenStudy (anonymous):

hmm that's how you solve a linear equation i suppose